OpenStudy (anonymous):

I need help with multi-step equations. The problem is 14n-8=34 Do I add 8 to each side or subtract?

OpenStudy (watchmath):

You want to make the n to stay by itself. So firs we need to get rid off the -8 by adding 8 to both sides 14n=34+8 14n=42 Next we want to get rid off the 14, so we divide by 14 n=42/14 n=3.